PartnerObserving Log for 2007 Mar 15/16

Observers: Rick Pogge, Peter Garnavich
Inst. Support: John Hill
TO: Ed Lundin

UTC Mar 15 2200: Day crew is replacing the spectacularly burned out brushes in the right-front elevation drive motor. When they are done we can start in on addressing the stuck shutter problem that re-emerged last night.

2300 - The telescope is now working with all 4 elevation drives and in balance. Starting work on the shutter. They're working on it on the scissors lift rather than taking it down to the lab. The IssueTrak report will detail what was done.

All subsequent times are UTC 2007 March 16

0120 - Back on the telescope, readying for tests.

0155 - Sky at sunset shows thick broken high clouds in the northern third of the sky, much of the thinner high clouds have broken up. Satellite shows a band of high cloud heading down from the north - conditions will not be photometric, and seeing might not be so hot, either.

0204 - Started some imaging tests while they reset the elevation drives after a runaway fault. Using skyflat test images and watching the logs. Doing a series of sky flats just to exercise the shutter and see if we're back in business. No telescope interaction at this phase, or tracking, so we'll not get dithers and stars will trail. Shutter is operational.

0215 - Sending scope to 7h/+27 for some sky flats in R and g (what we didn't get last night). Delay getting going because of some elevation drive faults that took a while to clear up. Didn't really get very far given some setup glitches. Given the amount of stuff done, I'm not surprised we're a little off at the start.

0225 - Tuning up primary mirror. Took our time as it was way out at the beginning. Also using this as a chance to learn more from John about the fine art of predicting corrections from pupil images, and thus informing us about how to read the lbcfpia output.

0301 - M67 B for first shot to assess seeing. Adjusting primary and then taking a snapshot to assess seeing at B. If good we'll proceed, otherwise we'll switch to a seeing-appropriate target. 5s B snap = 4pix FWHM = 0.9"

0312 - Started M67 B1, seeing 0.9", for project completion. First image has 4pix FWHM, so 0.9", proceeding..., 2nd image seeing a little better at 3.8pix FWHM, 3rd still looking good, 3.8pix, ... 8x180s images total in this sequence. Seeing got to 0.8" during 032829 (5th in sequence). Same in 6th. Nice images, none of the ugly PSFs we had last night in long exposures. All the stars are nice and round. Judging the seeing as between 0.8 and 0.9" during the first sequence. Satellite crossed Chip 3 and part of Chip 2 during the last image.

Satellite maps show the clouds from the north dissipating, but humidity is a little higher than last night (25%). Obscuration threat is diminishing, but still likely to remain nonphotometric.

0344 - Mirror tweak-up, then starting M67 B2. Minor change in the primary. First of 9 started at 0348. Lost guide star 40s into observation 2 of the sequence and never recovered. The seeing is 0.8-0.9" and we're in M67 in B for heaven's sake, how can you not find a suitable guide star. This autoguiding system is seriously broken. Image quality was degraded with a little bit of trailing. 3rd image picked up guide star OK and held it for the duration. Proceeding through the 9 images in this OB. Guide star started getting a little fatter during image 5 of the sequence - maybe 1" or so now. A couple of puffs in the guide star during image 6, seeing ~1". During image 7 the guide star was lost briefly about 130s into the exposure, but this time it recovered at about 142s elapsed time. Stars are not quite as round as previously. Image 8 was better, but still a little elongated with seeing ~0.9". Last image about same, seeing ~1", so conditions are not as good as during the first OB. We'll keep rolling.

0423 - Mirror tweak-up before starting M67 B3. A little astigmatism and coma, so could be part of the change in image quality at the end. Starting B3 at 0427, first image of 8 started at 0428 (takes about 1m for the mirror to move into place after the tweak-up). First image is much nicer, round and 0.8". 2nd and 3rd about the same. 0.8-0.9" seeing and round images for the most part. Image 6 was 3.5pix and nice and round, about the best image of the sequence thus far. Humidity up to 30%. Image 7 about as good as Image 6. Image 8 a little fuzzier and 3.8pix.

Finished the M67 project

0501 - Going to CosmosU1 field. Tweaking primary - Started cosmos_SDTUspec_1a at 0505.
  • Image 1: Lost 8s guide star part way through (didn't notice when), did not recover. Seeing ~1" and images out of round.
  • Image 2: Did not find a guide star at all. Looked at the technical chip images. They're pretty sparse, but the stars that are there should have been adequate to guide on with longer (16s) exposures, or even 8s exposures. Seeing ~1".
  • Image 3: started out w/o a guide star, but caught one when it ramped up to a 32s technical chip exposure. Not sure why it didn't ramp up the exposure times on images 1 and 2 to get guide stars. Image quaiity is much better - round and 1" seeing. Streak across chips 1 & 2 (fuzzy, not sure what it is).
  • Image 4: did not grab a guide star until about 40s into the exposure, then lost it about 20s later and did not recover. Images degraded (not round).
  • Image 5: got guide star at 8s then lost it 30s in and never recovered or attempted to ramp up the exptime. Seeing >1" and not very round.

The guide star algorithm is failing badly when there are not a lot of stars, which is a big problem in U at relatively high latitudes. We're going to stop observing this field and move on to good seeing targets in other bands where the good seeing won't be compromised by poor guider performance at U. We cannot achieve the seeing requirements for this project in the current state.

0541 - Going to Lockman Block 2 (V images) - using Vfastextra for mirror tweak, started at 0548 with Lockman VLA_V_1. A repeat from last night, when we had bad problems emerging.
  • Image 1: Locked on and held guide star the entire 360s exposure, ~1" seeing. Hard to find stars in this field, it is lousy with galaxies
  • Image 2: Locked on, star a little faint, so guiding may not be the best. Seeing now ~1-1.1"
  • Image 3: Good guide star locked, seeing ~1", images not as round as they could be.
  • Image 4: Guider lost star about 40s in (8s tech chip exptime). Tech chip 1 has a nice saturated star. Why can't it ramp exptime down sometimes? Images turned out lopsided.
  • Image 5: Guider lost star about 30s in (8s tech chip). Never recovered. Image took a very long time to appear on the Repository, a

0628 - Tweaked up primary, then executed Lockman VLA_V_2, the one that aborted last night when the elevation drive paniced.
  • Image 1: got a star, looks better than previously. Images are rounder, seeing 1" or better.
  • Image 2: got a star, also looks better. Images round and seeing ~1".
  • Image 3: failed to get guidestar lock w/8s. Poked about in the log analyzer, saw notation "IA System error [src/trackers/trackers.c:1879]" at about the time it lost lock and never recovered. Not sure if this means anything, but the error was absent before from the start of the exposure, and there were reports of "working with 2 guide stars".
  • Image 4: Guide star a bit fuzzier, seeing may be starting to degrade measureably. Wind appears to have picked up some (now 7m/s) The guide star was puffing and bounding around for a while there. Seeing on the final images was ~1.5" Conditions appear to be deteriorating.
  • Image 5: Got a star and locked, looks better but seeing is 1.1"

0706 - Setting for XMM_V_1, but will take a V snapshot after the primary tweak-up to see what the seeing is like. - Seeing is now >1" (1.2"), conditions have deteriorated past their stated seeing limits as judged from what they rejected in the past so we won't execute these OBs.

0719 - Going to Quasar/Point3, starting with _gr_3 OB. Primary tweak, then starting at 0723.
  • Image 1 (r) - seeing ~.9", better than before, but a little out of round
  • Image 2 (g) - it chose a galaxy as the guide star. Weird. Seeing ~1" but rounder than Image 1, streak across Chip 3
  • Image 3 (r) - Got a good star, seeing ~1", but not round.
  • Image 4 (g) - picked a star this time, not that round, seeing 1.1" (seeing is clearly variable)
  • Image 5 (r) - got a guide star, seeing ~1.1", not that round
  • Image 6 (g) - got a guide star, seeing 1" and round. Image quality is rather inconsistent (or our measurements aren't that robust)
  • Image 7 (r) - got a nice guide star, ~1" but not very round. Better on Chip 3 (round and ~0.9")
  • Image 8 (g) - good guide star, ~1" and round.

0738 - doing the _U_3 OB now (5x300s)
  • Image 1: has a guide star - images are nicely round and FWHM 4.7 pixels (1.07"). Nice.
  • Image 2: has a guide star, good lock, nice round images, 4.6pix (1.04"). Seeing has definitely settled down compared to the Lockman Hole fields.
  • Image 3: good star again, seeing now ~0.9" with round images. Things just got a lot better than when we started
  • Image 4: good star, not quite as round but good seeing (4pix FWHM)
  • Image 5: another nice star, round images and 3.8pix FWHM.

This finishes Quasar/Point3 Seeing was much better than when we started out. Xiaohui got a bonus. The U images are terrific.

0811 - going back to finish Lockman Block 2 since the conditions appear to have have improved, and the airmass window on it is still open. At least, we hope they have in that direction...First tweaking the primary, then starting XMM_V_1 at 0816, 5x360s:
  • Image 1: has a guide star, nice and round, 0.9" seeing - things really improved.
  • Image 2: looks like it's trying to guide on a galaxy, but the images are round and ~0.8" seeing.
  • Image 3: found a star this time (though this is the Lockman Hole, you can't be sure). Nice and round, 0.8" seeing.
  • Image 4: found a galaxy this time, stars are round and 0.8" seeing.
  • Image 5: star this time (or compact galaxy), images round, seeing ~0.75" - about the nicest images we've seen at V yet tonight.

0854 - checking primary, then running XMM_V_2 at 08:58
  • Image 1: got a guide star, seeing ~0.9", mostly round stars.
  • Image 2: got a guide star, much better, seeing ~0.75", round stars
  • Image 3: guiding on a galaxy, 0.8" seeing, round stars
  • Image 4: guiding on a galaxy, seeing more like 0.9" now, but stars are round
  • Image 5: guiding on a compact galaxy, round stars, a little over 0.8" seeing

Completed Lockman Block 2, though the image quality for the first set (VLA_V_x) needs to be checked by the PI.

0945 - Attempting Zaritsky/NGC4826 U observation (fits in a schedule hole given seeing/airmass/non-photometric), tweaked up primary. First image came in a little out of focus. Aborted OB during 2nd exposure and re-checked the focus. Offsett 10arcmin from current point (a bright star got grabbed and gave garbage with lbcfpia) and doing Bfastextra again. Got the same result as near the galaxy. Took a snapshot and got crappy images. Weird. Mirror actuators all look correct. Not sure what the deal is, but this is pretty hopeless

Wind has gone down and it is warming up, Temp=7C. More like a summer night, not March.

1020 - Trying another Zaritsky field, NGC4244, still acting perplexing. Seeing 1.5". That's no good. Going to go do something that can tolerate poor seeing.

1043 - trying Dwarf/fuobj11, the one aborted last night by the shutter failure. Started fuobj11_1_r,
  • Image 1: lost the guide star within 60s of start..., never recovered. Stars turned out reasonably round, but seeing 1.7"
  • Image 2: GUI locked up during start of the 2nd image of the 4 but it is still taking data. Got its attention back by changing pages, but it took a while for it to service the request. 2nd image had a guide star, seeing more like 1.4-1.5" now.
  • Image 3: Lost guide star mid-way through the 300s exposure, never recovered. GUI locked up again, now 3rd images hasn't made it to the Repository yet. Did't appear until #4 below, seeing ~1.3-1.4"
  • Image 4: Guide star looks smaller, maybe seeing is settling down. Seeing ~1.2"

1113 - checking Primary, then running fuobj11_2_g @ 1116.
  • Image 1: got a nice guide star, a little out of round, seeing a little over 1"
  • Image 2: star bounced around a bit, and fuzzier at times. 1.1" seeing, not that round
  • Image 3: Star is fuzzy and varying, seeing getting variable, but still turned out 1.1", though not as round as it could be
  • Image 4: Star found, a little smaller this time, seeing still ~1.1" and slightly out of round.

1139 - checking Primary, then running fuobj11_2_r @ 1142
  • Image 1: compact guide star, good guiding, ~0.8" seeing and roundness OK, nice image.
  • Image 2: good guide star, seeing ~0.9", reasonably round
  • Image 3: guide star fuzzy (galaxy)?, seeing ~0.9"
  • Image 4: compact guide star, seeing ~0.9", mostly round.

Dwarf/fuobj11 completed remaining 3 OBs

1208 - doing 25bias and 900s dark, which we didn't get in this afternoon, while waiting for astronomical twilight for sky flats.

1245 - Scope at 14h/+27 for twilight flats. V, g, and r then U.
  • SkyFlatV_5 scaled 10x, range 9000-20000,
  • SkyFlatG_5 scaled 3x, range 17K-30K,
  • SkyFlatR_5 scaled 1x, range 7K-9K for 2, paused then 2x scaled for 2 (125827 & 125906), range 24K-29K, scaled 1.0x for 4 (130100 - 130253), range 24K - 39K.
  • SkyFlatU_5 scaled 1x, range 20K-40K

That's it, we're done.

-- RichardPogge - 15 Mar 2007
Topic revision: r10 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback